# Feature-aware Hypergraph Generation via Next-Scale Prediction

### v1

@@ Abstract
Hypergraphs generalize traditional graphs by allowing hyperedges to connect multiple nodes, making them well-suited for modeling complex structures with higher-order relationships, such as 3D meshes, molecular systems, and electronic circuits. While topology is central to hypergraph structure, many real-world applications also require node and hyperedge features. Existing hypergraph generation methods focus solely on topology, often overlooking feature modeling. In this work, we introduce FAHNES (feature-aware hypergraph generation via next-scale prediction), a hierarchical approach that jointly generates hypergraph topology and features. FAHNES builds a multi-scale representation through node coarsening, then learns to reconstruct finer levels via localized expansion and refinement, guided by a new node budget mechanism that controls cluster splitting. We evaluate FAHNES on synthetic hypergraphs, 3D meshes, and molecular datasets. FAHNES achieves competitive results in reconstructing topology and features, establishing a foundation for future research in featured hypergraph generative modeling.
@@ 1 Introduction
Hypergraphs extend traditional graphs by allowing edges—called hyperedges—to connect more than two nodes, enabling a natural representation of complex, multi-way interactions. This expressiveness makes hypergraphs well-suited for modeling systems in molecular chemistry, electronic circuit design, and 3D geometry. As a result, they have been applied in diverse areas such as drug discovery [1], contagion modeling [2], recommendation systems [3], molecular biology [4], and urban planning [5].
Despite their wide applicability, generative models for hypergraphs remain relatively underexplored. Existing approaches primarily focus on generating topology alone [6], neglecting the generation of node and hyperedge features. As shown in Figure 1, the sequential generation of topology followed by features is ineffective. Joint generation is particularly challenging for hypergraphs due to their variable-size hyperedges and higher structural complexity. Moreover, existing methods for featured graph generation use flat, non-hierarchical architectures, which scale poorly [7, 8, 9].
In this work, we introduce a feature-aware hypergraph generation via next-scale prediction (FAHNES) model, a novel hierarchical approach for jointly generating hypergraph topology and features. Inspired by multi-scale generative strategies in image synthesis [10], FAHNES leverages a coarsening-expansion strategy [6, 11]: during training, it merges nodes based on connectivity to build a hierarchical representation, then learns to progressively recover finer-scale hypergraphs. A key contribution is the use of a budget mechanism that tracks the number of permissible expansions per cluster. This enables global structural constraints to influence local generation, helping to mitigate the limited receptive fields of deep models. Additionally, we adapt flow-matching optimal transport (OT) coupling techniques [12, 13] to learn straighter and more consistent flows during generation. Our main contributions are:
  • We introduce the first hierarchical model for joint generation of hypergraph topology and features targeting specific data distributions, such as 3D meshes and molecules.
  • We propose a novel budget-based mechanism that improves global consistency during generation for hierarchical methods.
  • We generalize minibatch OT-coupling to our hypergraph generative setting.
  • We validate our approach on both synthetic and real-world datasets, including molecular structures and 3D meshes, showing promising results for hierarchical generative methods.
@@ 2 Related Work
Graph and hypergraph generation using deep learning. Graph generation has seen significant advances in recent years. Early approaches, such as GraphVAE [14], employed autoencoders to embed graphs into latent spaces for sampling. Subsequent models leveraged recurrent neural networks to sequentially generate adjacency matrices, improving structural fidelity [15]. More recently, diffusion-based methods have enabled permutation-invariant graph generation [16, 7], with extensions incorporating structural priors such as node degrees [17]. Many of these previous methods jointly model node features and topology, but are limited to small graphs due to scalability challenges. For example, diffusion-based models [7, 8, 9] operate over entire graphs by gradually corrupting structures and features, then training models to denoise them. However, their scalability is constrained by the combinatorial nature of edge prediction in large graphs.
@@
To mitigate this, hierarchical methods have been proposed. For example, Bergmeister et al. [11] introduced a scalable graph generation framework based on a coarsen-then-expand approach, merging nodes to form coarse representations and progressively reconstructing finer details. This framework was extended to hypergraphs by Gailhard et al. [6], which allows edges to connect more than two nodes. However, both methods focus exclusively on topology generation, neglecting node and hyperedge features essential for many applications.
@@
Applications of hypergraph generation. Generative models that capture both higher-order structure and node/hyperedge features are critical in numerous domains. In molecular design, regular graph generative models struggle to accurately represent rings and scaffolds [7], which naturally correspond to hyperedges involving multi-atom interactions rather than pairwise bonds. Similarly, 3D shape modeling involves surfaces such as triangles, quads, or general polygons that extend beyond simple pairwise connectivity. Conventional approaches often rely on fixed topology, quantization, or autoregressive sequence modeling with transformers [18, 19], limiting their flexibility and scalability. Hypergraphs provide a more expressive framework by treating faces as hyperedges, enabling the joint generation of topology and features.
In contrast to previous approaches that either target hypergraph topology alone or feature-aware graph generation, we introduce the first method that jointly models both hypergraph structure and features within a unified and scalable framework. FAHNES extends hierarchical generative methods with feature-aware expansions, a novel node budget mechanism, and flow-matching training adapted to our strategy. This combination enables effective generative modeling for featured hypergraphs.
@@
3 Feature-aware Hypergraph Generation via Next-scale Prediction
((a)) Hypergraph
((b)) Clique expansion
((c)) Star expansion
Figure 2: Comparison of a hypergraph and its clique and star expansions.
[…]
@@ 3 Feature-aware Hypergraph Generation via Next-scale Prediction
Basic definitions. We define a graph $G=(\mathcal{V},\mathcal{E})$ as a pair consisting of a set of vertices $\mathcal{V}$ and a set of edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Graphs may also carry node and edge features, represented by matrices $\mathbf{F}_{\mathcal{V}}\in\mathbb{R}^{|\mathcal{V}|\times m}$ and $\mathbf{F}_{\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|\times l}$, where $m$ and $l$ denote the dimensionality of the features. Each edge $e\in\mathcal{E}$ corresponds to a pair $(u,v)$, indicating a connection between nodes $u$ and $v$. A bipartite graph $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$ is a special case where the vertex set is split into two disjoint subsets $\mathcal{V}_{L}$ and $\mathcal{V}_{R}$, and edges exist only between the two parts, i.e., $\mathcal{E}\subseteq\mathcal{V}_{L}\times\mathcal{V}_{R}$. The full set of nodes is $\mathcal{V}=\mathcal{V}_{L}\cup\mathcal{V}_{R}$. In this work, we consider node features for bipartite graphs, denoted by $\mathbf{F}_{L}$ for the left-side nodes and $\mathbf{F}_{R}$ for the right-side nodes.
@@
A hypergraph $H$ extends the concept of a graph and is specified by a pair $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the vertex set and $\mathcal{E}$ comprises hyperedges, with each $e\in\mathcal{E}$ being a subset of $\mathcal{V}$. The key distinction of hypergraphs lies in their ability to connect arbitrary numbers of vertices through a single hyperedge. Similar to graphs, hypergraphs can possess node and edge features, which are denoted $\mathbf{F}_{\mathcal{V}}$ and $\mathbf{F}_{\mathcal{E}}$. We consider two fundamental graph-based representations of hypergraphs: clique and star expansions. The clique expansion transforms a hypergraph $H$ into a graph $C=(\mathcal{V}_{c},\mathcal{E}_{c})$, where $\mathcal{E}_{c}=\{(u,v)\mid\exists~{}e\in\mathcal{E}:u,v\in e\}$. The star expansion converts a hypergraph $H$ into a bipartite structure $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E}_{b})$, where $\mathcal{V}_{L}=\mathcal{V}$, $\mathcal{V}_{R}=\mathcal{E}$, and $\mathcal{E}_{b}=\{(v,e)\mid v\in\mathcal{V}_{L},e\in\mathcal{V}_{R},v\in e\text{ in }H\}$. Left side nodes $\mathcal{V}_{L}$ represent the nodes of the hypergraph, while right side nodes $\mathcal{V}_{R}$ represent hyperedges. We provide a visual example of the clique and star expansions in Figure 2.
Our work aims to develop a model capable of sampling from the underlying distribution of a given featured hypergraph dataset $(H_{1},\dots,H_{N})$, i.e., learning the joint probability of features and topology of this distribution of hypergraphs. For detailed mathematical proofs of all propositions and lemmas presented in this work, please consult the supplementary material.
@@ 3.1 Overview
Previous work [6] simplifies hypergraph generation by reducing it to standard graph generation through two projections: i) the clique expansion where every hyperedge becomes a clique that links all its incident nodes, and ii) the star expansion, which produces a bipartite graph by introducing one node per hyperedge and connecting it to the nodes it contains. While conceptually elegant, this approach faces significant challenges: i) it struggles to produce the correct number of nodes, especially in large hypergraphs, and ii) it does not support the generation of node and hyperedge features, limiting its applicability to real-world, featured domains.
To overcome the limitations of prior approaches, we propose two key contributions. First, we introduce a node budget mechanism, in which generation begins with a single super-node carrying the full node budget. At each generation step, this budget is recursively split among child nodes. This approach enables a more localized and structured control over the final node count, compared to previous methods that simply concatenate the desired size to node embeddings. Second, we propose hierarchical feature generation, where each expansion step predicts the mean feature of its future child nodes, conditioned on the parent’s feature. This hierarchical feature generation scheme is thus a generalization of [20] to hypergraphs, where the prediction at current scale conditions the predictions at next scale. To train FAHNES, we adopt the flow-matching framework [21], where the generation process is formulated as learning to reverse a known noising procedure. Specifically, the model is trained to recover the true values for node and hyperedge expansions, edge deletions, budget allocations, and features, all of which are subjected to noise. Our workflow is illustrated in Figure 3.
Figure 3: Our framework adopts a coarsening-expansion strategy. i) During training, input hypergraphs are progressively coarsened by merging nodes and hyperedges, yielding a multiscale representation. Node features are averaged during merging, and budgets are summed. ii) The model is trained to predict which nodes were merged at each scale. iii) In the expansion phase, merged nodes are expanded (shown in dark), inheriting their parent’s features, budget, and connectivity. The model is trained to (a) identify which edges should be removed (dotted lines), (b) predict how the parent’s budget should be split across the children (instead of predicting absolute values), and (c) refine the features of newly expanded nodes (refinement).
@@ 3.2 Budgeted Coarsening
During coarsening, we keep track of two quantities for every cluster we create. First, budgets count how many nodes the cluster already contains. Every node and hyperedge begins with a budget of $1$; when several of them merge, their budgets are summed and assigned to the new super-node or super-hyperedge. Secondly, features summarize the attributes of the merged vertices. We use a weighted average: the feature of a cluster is the budget-weighted mean of the features of its members.
@@
Setting cluster features as the average of the nodes they contain minimizes the mean squared error (MSE) between the features of the fully expanded hypergraph and those of the original hypergraph.
In what follows we denote left-side (node) budgets and features by $\mathbf{b}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ and $\mathbf{F}_{L}$, and right-side (hyperedge) budgets and features by $\mathbf{b}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$ and $\mathbf{F}_{R}$.
Definition 2 (Bipartite graph coarsening).
Let $H$ be an arbitrary hypergraph, $C=(\mathcal{V}_{c},\mathcal{E}_{c})$ its unfeatured weighted clique expansion, and $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E},\mathbf{b}_{L},\mathbf{b}_{R},\mathbf{F}_{L},\mathbf{F}_{R})$ its featured bipartite representation. Given a partitioning $\mathcal{P}_{L}=\{\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(n)}\}$ of the node set $\mathcal{V}_{L}$ such that each part $\mathcal{V}^{(p)}$ induces a connected subgraph in $C$, we construct an intermediate coarsening $\tilde{B}(B,\mathcal{P}_{L})=(\bar{\mathcal{V}}_{L},\mathcal{V}_{R},\bar{\mathcal{E}},\bar{\mathbf{b}},\bar{\mathbf{F}}_{L},\mathbf{F}_{R})$ by merging each part $\mathcal{V}^{(p)}$ into a single node $v^{(p)}\in\bar{\mathcal{V}}_{L}$, and by defining:
$\left(\bar{\mathbf{b}}_{L}\right)_{v^{(p)}}=\sum_{v\in\mathcal{V}^{(p)}}\left(\mathbf{b}_{L}\right)_{v},\qquad\left(\bar{\mathbf{F}}_{L}\right)_{v^{(p)}}=\frac{1}{\left(\bar{\mathbf{b}}_{L}\right)_{v^{(p)}}}\sum_{v\in\mathcal{V}^{(p)}}\left(\mathbf{b}_{L}\right)_{v}\left(\mathbf{F}_{L}\right)_{v},$ (1)
for every supernode $\mathcal{V}^{(p)}$. An edge $e_{\{p,q\}}\in\bar{\mathcal{E}}$ is added between $v^{(p)}\in\bar{\mathcal{V}}_{L}$ and $v^{(q)}\in\mathcal{V}_{R}$ if there exists an edge $e_{\{i,q\}}\in\mathcal{E}$ in the original bipartite representation between some $v^{(i)}\in\mathcal{V}^{(p)}$ and $v^{(q)}$.
To complete the coarsening process, we define an equivalence relation $v_{1}\sim v_{2}\iff\mathcal{N}(v_{1})=\mathcal{N}(v_{2})$ on $\mathcal{V}_{R}$, where $\mathcal{N}(v)$ denotes the set of neighbors of $v$. This induces a partitioning $\mathcal{P}_{R}=\{\mathcal{V}_{R}^{(1)},\dots,\mathcal{V}_{R}^{(m)}\}$, allowing us to construct the fully coarsened bipartite representation $\bar{B}(\tilde{B},\mathcal{P}_{L})=(\bar{\mathcal{V}}_{L},\bar{\mathcal{V}}_{R},\bar{\mathcal{E}},\bar{\mathbf{b}},\bar{\mathbf{F}}_{L},\bar{\mathbf{F}}_{R})$ by merging each part $\mathcal{V}_{R}^{(p)}$ into a single node $v_{R}^{(p)}\in\bar{\mathcal{V}}_{R}$, similarly to the construction of $\bar{\mathcal{V}}_{L}$, and by defining:
$\left(\bar{\mathbf{b}}_{R}\right)_{v^{(p)}}=\sum_{v\in\mathcal{V}_{R}^{(p)}}\left(\mathbf{b}_{R}\right)_{v},\qquad\left(\bar{\mathbf{F}}_{R}\right)_{v^{(p)}}=\frac{1}{\left(\bar{\mathbf{b}}_{R}\right)_{v^{(p)}}}\sum_{v\in\mathcal{V}_{R}^{(p)}}\left(\mathbf{b}_{R}\right)_{v}\left(\mathbf{F}_{R}\right)_{v},$ (2)
for every superhyperedge $\mathcal{V}_{R}^{(p)}$.
Remark 3.
Informally, we first pick the node clusters on the clique expansion, merge those nodes on the left side of the bipartite graph, and compute their budgets and features as above. We then merge right-side nodes that appear multiple times, i.e., they connect the same left-side nodes. In practice, for the final reduction step, when only one node and one hyperedge remain, we also replace their features by matrices full of zeros. This design choice enables a fully agnostic initial step during generation.

### v2

# Feature-aware (Hyper)graph Generation via Next-Scale Prediction

@@ Abstract
Graph generative models have shown strong results in molecular design but struggle to scale to large, complex structures. While hierarchical methods improve scalability, they usually ignore node and edge features, which are critical in real-world applications. This issue is amplified in hypergraphs, where hyperedges capture higher-order relationships among multiple nodes. Despite their importance in domains such as 3D geometry, molecular systems, and circuit design, existing generative models rarely support both hypergraphs and feature generation at scale. In this paper, we introduce FAHNES (feature-aware hypergraph generation via next-scale prediction), a hierarchical framework that jointly generates hypergraph topology and features. FAHNES builds multi-scale representations through node coarsening and refines them via localized expansion, guided by a novel node budget mechanism that controls granularity and ensures consistency across scales. Experiments on synthetic, 3D mesh and graph point cloud datasets show that FAHNES achieves state-of-the-art performance in jointly generating features and structure, advancing scalable hypergraph and graph generation.
@@ 1 Introduction
Generating discrete geometric structures such as graphs and hypergraphs is an important challenge in modern machine learning [1]. These structures are central to applications ranging from molecular design and materials discovery to electronic circuits and 3D shape modeling [2, 3, 4, 5]. Their ability to capture complex relationships—pairwise in graphs, and multi-way in hypergraphs—makes them an essential tool for representing and synthesizing structured data. In many real-world settings, topology alone is not enough: node and edge (or hyperedge) features often carry important semantic or geometric information, such as atom and bond types in molecules, or vertex coordinates in 3D meshes.
@@
Despite recent advances, existing methods for featured graph generation struggle to scale. Most of these approaches use flat architectures that model the entire structure at once, leading to quadratic computational and memory complexities [6, 7, 8]. This limits their use to moderately sized graphs. To address scalability, hierarchical generation strategies have been proposed for both graphs and hypergraphs [9, 10]. These methods build the structure in stages, starting from a coarse representation and progressively upsampling and refining it, which allows them to handle much larger topologies.
However, existing hierarchical methods only focus on unfeatured structures [9, 10]. Extending them to generate features is challenging because different regions of a graph or hypergraph often grow at uneven rates during the refinement process, making it difficult to maintain consistency across scales. Furthermore, sequentially generating topology first and then features—rather than modeling them jointly—is ineffective in complex settings, as illustrated in Figure 1.
To overcome these limitations, we propose FAHNES  (feature-aware hypergraph generation via next-scale prediction), a hierarchical generative framework that jointly models topology and features for both graphs and hypergraphs [^1] Since hypergraphs are a generalization of graphs, we present FAHNES on hypergraphs. The explanation for graphs follows straightforwardly.. FAHNES builds multi-scale representations [11] through node coarsening and reconstructs fine structures via localized expansion, while directly predicting features alongside structure at each stage. A fundamental component of our model is a node budget mechanism, which encodes local growth constraints and helps maintain consistency in regions that expand at different rates. In addition, we adapt minibatch flow-matching optimal-transport (OT) coupling [12, 13] to the hierarchical setting, improving stability and coherence during generation. This combination enables scalable generation of large, featured graphs or hypergraphs across diverse domains, from 3D meshes to point clouds. Our main contributions are:
  • We introduce the first scalable hierarchical model for the joint generation of hypergraph topology and features, targeting complex data such as 3D meshes and point clouds.
  • We propose a novel budget-based mechanism that enhances global structural coherence in hierarchical generation.
  • We extend minibatch OT-coupling to hierarchical graph or hypergraph generation, improving generative stability and quality.
  • We validate FAHNES on both synthetic and real-world datasets, demonstrating strong performance in jointly generating topology and features.
@@ 2 Related Work
Graph and hypergraph generation using deep learning. Graph generation has seen significant advances in recent years. Early approaches, such as GraphVAE [14], employed autoencoders to embed graphs into latent spaces for sampling. Subsequent models leveraged recurrent neural networks to sequentially generate adjacency matrices, improving structural fidelity [15]. More recently, diffusion-based methods have enabled permutation-invariant graph generation [16, 6], with extensions incorporating structural priors such as node degrees [17]. Many of these previous methods jointly model node features and topology, but are limited to small graphs due to scalability challenges. Diffusion-based models [6, 7, 8] operate over complete graphs by gradually corrupting structures and features, then training models to denoise them. However, their scalability is constrained by the combinatorial number of possible edges in graphs.
@@
To mitigate this, hierarchical methods have been proposed. [9] introduced a scalable graph generation framework based on a coarsen-then-expand approach, merging nodes to form coarse representations and progressively reconstructing finer details. This framework was extended to hypergraphs by [10], which allows edges to connect more than two nodes. However, both methods focus exclusively on topology generation, neglecting node and hyperedge features essential for many applications.
@@
Applications of hypergraph generation. Generative models that capture both higher-order structure and node/hyperedge features are critical in numerous domains. In molecular design, regular graph generative models struggle to accurately represent rings and scaffolds [6, 18], which naturally correspond to hyperedges involving multi-atom interactions rather than pairwise bonds. Similarly, 3D shape modeling involves surfaces such as triangles, quads, or general polygons that extend beyond simple pairwise connectivity. Conventional approaches often rely on fixed topology, quantization, or autoregressive sequence modeling with transformers [19, 20], limiting their flexibility and scalability. Hypergraphs provide a more general framework by treating faces as hyperedges, enabling the joint generation of topology and features.
Unlike prior work that focuses solely on topology or relies on flat, non-scalable designs, we present the first unified, scalable approach for jointly modeling hypergraph structure and features. FAHNES enhances hierarchical generative modeling with feature integration, a novel node-budget mechanism, and flow-matching training tailored to our strategy.
@@
3 Feature-aware (Hyper)graph Generation via Next‑scale Prediction
Notations. Throughout this paper, we use calligraphic letters, like $\mathcal{V}$, to represent sets, with their cardinality denoted by $|\mathcal{V}|$. Matrices are represented by bold uppercase letters (e.g., $\mathbf{A}$), while vectors are indicated by bold lowercase letters (e.g., $\mathbf{x}$). The transpose and point-wise multiplication operations are denoted by $(\cdot)^{\top}$ and $\odot$, respectively. $\lceil\cdot\rfloor$ denotes rounding to the nearest integer.
@@ 3 Feature-aware (Hyper)graph Generation via Next‑scale Prediction
Basic definitions. We define a graph $G=(\mathcal{V},\mathcal{E})$ as a pair consisting of a set of vertices $\mathcal{V}$ and a set of edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$. Graphs may also carry node and edge features, represented by matrices $\mathbf{F}_{\mathcal{V}}\in\mathbb{R}^{|\mathcal{V}|\times m}$ and $\mathbf{F}_{\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|\times l}$, where $m$ and $l$ denote the dimensionality of the features. Each edge $e\in\mathcal{E}$ corresponds to a pair $(u,v)$, indicating a connection between nodes $u$ and $v$. A bipartite graph $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E})$ is a special case of a graph where the vertex set is split into two disjoint subsets $\mathcal{V}_{L}$ and $\mathcal{V}_{R}$, and edges exist only between the two parts, i.e., $\mathcal{E}\subseteq\mathcal{V}_{L}\times\mathcal{V}_{R}$. The full set of nodes is $\mathcal{V}=\mathcal{V}_{L}\cup\mathcal{V}_{R}$. In this work, we consider node features for bipartite graphs, denoted by $\mathbf{F}_{L}$ for the left-side nodes and $\mathbf{F}_{R}$ for the right-side nodes.
@@
A hypergraph $H$ extends the concept of a graph and is specified by a pair $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the vertex set and $\mathcal{E}$ comprises hyperedges, with each $e\in\mathcal{E}$ being a subset of $\mathcal{V}$. The main distinction of hypergraphs lies in their ability to connect arbitrary numbers of vertices through a single hyperedge. Similar to graphs, hypergraphs can possess node and hyperedge features, which are denoted $\mathbf{F}_{\mathcal{V}}$ and $\mathbf{F}_{\mathcal{E}}$. We consider two fundamental graph-based representations of hypergraphs: clique and star expansions. The clique expansion transforms a hypergraph $H$ into a graph $C=(\mathcal{V},\mathcal{E}_{c})$, where hyperedges are replaced by cliques, i.e., $\mathcal{E}_{c}=\{(u,v)\mid\exists~e\in\mathcal{E}:u,v\in e\}$. The star expansion converts a hypergraph $H$ into a bipartite structure $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E}_{b})$, where $\mathcal{V}_{L}=\mathcal{V}$, $\mathcal{V}_{R}=\mathcal{E}$, and $\mathcal{E}_{b}=\{(v,e)\mid v\in\mathcal{V}_{L},e\in\mathcal{V}_{R},v\in e\text{ in }H\}$. Left side nodes $\mathcal{V}_{L}$ represent the nodes of the hypergraph, while right side nodes $\mathcal{V}_{R}$ represent hyperedges.
Our objective is to develop a generative model that can sample from the underlying distribution of a dataset of featured hypergraphs (or graphs) $(H_{1},\dots,H_{N})$, i.e., to learn the joint distribution over both topology and features. We present FAHNES in the more general setting of hypergraphs; since graphs are a special case of hypergraphs, the model can be directly adapted to graphs, as discussed in Section 3.6. Complete mathematical proofs of all propositions in this paper appear in Appendix A.
@@ 3.1 Overview
FAHNES follows a hierarchical pipeline as shown in Figure 2: i) during training, multi-scale representations of input (hyper)graphs are produced; ii) a model learns to reconstruct a higher scale from a lower one. Our approach adopts a coarsening–expansion framework [10]. We first perform standard spectrum-preserving coarsening [21] on the clique expansion of the hypergraph to obtain the multi-scale representation. The process is then reversed through expansion and refinement, reconstructing both topology and features at progressively finer scales. Expansion and refinement are learned on the bipartite representation using the flow-matching framework [22], treating generation as the inverse of the coarsening process.
Figure 2: Our framework adopts a coarsening-expansion strategy. i) During training, input hypergraphs are progressively coarsened by merging nodes and hyperedges, yielding a multiscale representation. Node features are averaged during merging, and budgets are summed. ii) The model is then trained to predict which nodes were merged at each scale. iii) In the expansion phase, merged nodes (shown in dark in the leftmost column) are expanded back (copies shown in dark), inheriting their parent’s features, budget, and connectivity. In the refinement phase, the model is trained to (a) identify which edges should be removed (dotted lines), (b) predict how the parent’s budget should be split across the children, and (c) refine the features of newly expanded nodes.
To address uneven growth across regions, we introduce a node budget mechanism: each cluster is assigned a budget indicating its remaining expansions. Budgets are recursively divided among child nodes, starting from a single super-node with the full budget and ending when all clusters have a budget of one. This provides more local control over the final node count and improves global consistency compared to prior methods [9, 10], which only append the desired size to node embeddings. At each expansion step, the model also predicts the mean feature of future child nodes conditioned on the parent’s feature, extending the scale-wise autoregressive idea of [23] so that predictions at one scale guide those at the next.
@@ 3.2 Budgeted Coarsening
We adopt the coarsening strategy of [10], representing the hypergraph at multiple resolution levels by merging nodes into clusters. Node pairs to be merged are selected using the spectrum-preserving coarsening of [21], applied to the hypergraph’s clique expansion. Mergings in the clique expansion are then mirrored in the bipartite representation, with corresponding left side nodes being merged, and copies of the same hyperedge being subsequently collapsed. We limit node mergings to at most two per cluster, which consequently bounds hyperedge mergings to at most three [10]. For each resulting cluster, we track two quantities. First, the budget is the total number of nodes contained in the cluster. Initially, each node and hyperedge has a budget of $1$; when clusters merge, their budgets are summed and assigned to the new super-cluster. The node budget provides a local growth constraint, indicating how many fine-level nodes each coarse cluster should expand into and guiding refinement in regions that grow at different rates. Second, the super-cluster’s feature is computed as the weighted mean of the features of the clusters it absorbs, so it represents the average feature of all its contained nodes.
We denote node (left‑side) budgets and features by $\mathbf{b}_{L}\in\mathbb{N}^{|\mathcal{V}_{L}|}$ and $\mathbf{F}_{L}$, and hyperedge (right‑side) budgets and features by $\mathbf{b}_{R}\in\mathbb{N}^{|\mathcal{V}_{R}|}$ and $\mathbf{F}_{R}$. The following definition formalizes our coarsening process.
Definition 1 (Bipartite graph coarsening).
Let $H$ be an arbitrary hypergraph, $C=(\mathcal{V}_{c},\mathcal{E}_{c})$ its clique expansion, and $B=(\mathcal{V}_{L},\mathcal{V}_{R},\mathcal{E}_{b},\mathbf{b}_{L},\mathbf{b}_{R},\mathbf{F}_{L},\mathbf{F}_{R})$ its featured bipartite representation. Let $\mathcal{P}_{L}=\{\mathcal{V}^{(1)},\ldots,\mathcal{V}^{(n)}\}$ be a partitioning [^2] That is, $\mathcal{V}^{(p)}\subseteq\mathcal{V}_{L}$, $\bigcup_{i=1}^{{n}}\mathcal{V}^{(i)}=\mathcal{V}_{L}$, and $\mathcal{V}^{(i)}\cap\mathcal{V}^{(j)}=\emptyset~\forall~1\leq i,j\leq n$}. of the node set $\mathcal{V}_{L}$ such that each set $\mathcal{V}^{(p)}$ induces a connected subgraph in $C$. We construct an intermediate coarsening $\tilde{B}(B,\mathcal{P}_{L})=(\bar{\mathcal{V}}_{L},\mathcal{V}_{R},\bar{\mathcal{E}}_{b},\bar{\mathbf{b}}_{L},\mathbf{b}_{R},\bar{\mathbf{F}}_{L},\mathbf{F}_{R})$ by merging each part $\mathcal{V}^{(p)}$ into a single node $v^{(p)}\in\bar{\mathcal{V}}_{L}$, and by defining:
$\bar{\mathbf{b}}_{L}[p]=\sum_{v\in\mathcal{V}^{(p)}}\mathbf{b}_{L}[v],\qquad\bar{\mathbf{F}}_{L}[p]=\frac{1}{\bar{\mathbf{b}}_{L}[p]}\sum_{v\in\mathcal{V}^{(p)}}\mathbf{b}_{L}[v]\mathbf{F}_{L}[v],$ (1)
for every merged node $\mathcal{V}^{(p)}$. An edge $e_{\{p,q\}}\in\bar{\mathcal{E}}_{b}$ is added between $v^{(p)}\in\bar{\mathcal{V}}_{L}$ and $v^{(q)}\in\mathcal{V}_{R}$ if there exists an edge $e_{\{i,q\}}\in\mathcal{E}_{b}$ in the original bipartite representation between some $v^{(i)}\in\mathcal{V}^{(p)}$ and $v^{(q)}$.
To complete the coarsening process, we define an equivalence relation $v_{1}\sim v_{2}\iff\mathcal{N}(v_{1})=\mathcal{N}(v_{2})$ on $\mathcal{V}_{R}$, where $\mathcal{N}(v)$ denotes the set of neighbors of $v$, i.e., we consider right side nodes having the same set of neighbors as equivalent, or in other words we consider hyperedges containing the same set of nodes as equivalent. This induces a partitioning $\mathcal{P}_{R}=\{\mathcal{V}_{R}^{(1)},\dots,\mathcal{V}_{R}^{(m)}\}$, allowing us to construct the fully coarsened bipartite representation $\bar{B}(\tilde{B},\mathcal{P}_{L})=(\bar{\mathcal{V}}_{L},\bar{\mathcal{V}}_{R},\bar{\mathcal{E}}_{b},\bar{\mathbf{b}}_{L},\bar{\mathbf{b}}_{R},\bar{\mathbf{F}}_{L},\bar{\mathbf{F}}_{R})$ by merging each part $\mathcal{V}_{R}^{(p)}$ into a single node $v_{R}^{(p)}\in\bar{\mathcal{V}}_{R}$, similarly to the construction of $\bar{\mathcal{V}}_{L}$, and by defining:
$\bar{\mathbf{b}}_{R}[p]=\sum_{v\in\mathcal{V}_{R}^{(p)}}\mathbf{b}_{R}[v],\qquad\bar{\mathbf{F}}_{R}[p]=\frac{1}{\bar{\mathbf{b}}_{R}[p]}\sum_{v\in\mathcal{V}_{R}^{(p)}}\mathbf{b}_{R}[v]\mathbf{F}_{R}[v],$ (2)
for every merged hyperedge $\mathcal{V}_{R}^{(p)}$.
Remark 2.
Informally, we cluster nodes in the clique expansion, merge the corresponding left-side nodes of the bipartite graph, and compute their budgets and features as described above. Right-side nodes (hyperedges) connected to the same left-side nodes are then merged. In our implementation, we select nodes for clustering using spectrum-preserving coarsening [21]. In the final reduction—when only one node and one hyperedge remain—we replace their features with zero matrices to yield a feature-agnostic initialization for generation. Please note that cluster size vectors introduced in [9] correspond to the sizes of all sets $\mathcal{V}_{L}^{(p)}$ and $\mathcal{V}_{R}^{(p)}$, i.e., the number of mergings for each cluster at this specific step, and not the number of nodes absorbed since the initialization of coarsening—which we call budget. Further details about the coarsening methodology are provided in Appendix B. Some visual examples of coarsening sequences are provided in Appendix C.
We provide the following theoretical result regarding the feature merging process in FAHNES.
@@
Setting cluster features as the mean of the nodes they contain minimizes the mean squared error between the features of the fully expanded hypergraph and those of the original hypergraph.